how many atoms are there in 25g of sodium carbonate?
1
Expert's answer
2020-09-30T07:49:53-0400

formula of sodium carbonate is Na2CO3 and molar mass of it = (23*2)+(12*1)+(16*3) = 46+12+48 = 106 g\mol (here '*' corresponds to multiplication)

so mol of sodium carbonate = (mass of carbonate \ molar mass)

=(25\106) = 0.236 mol

now in one molecule of sodium carbonate contains 2 Na , 1 carbon and 3 oxygen atom , so total no of atom in one molecule is (2+1+3) = 6 no of atoms

we know from mole concept that 1 mol sodium carbonate contains avogadro no of sodium carbonate molecule that is = 6.023 * 1023 no of molecules = 6*6.023*1023 no of atoms.

so the no of atoms in 0.236 mol of sodium carbonate = 0.23*6*6.023*1023

= 8.311*1023 no of atoms. (ans)